When 25.8 mL of 0.500 M H2SO4 is added to 25.8 mL of 1.00 M KOH in a coffee-cup calorimeter at 23.50°C, the temperature rises to 30.17°C. Calculate ΔH of this reaction. (Assume that the total volume is the sum of the individual volumes and that the density and specific heat capacity of the solution are the same as for pure water.) (d for water = 1.00 g/mL; c for water = 4.184 J/g·°C.)

1
Expert's answer
2021-03-08T06:10:30-0500

"\u0394H= m\u00d7c\u00d7dT"

dT= 30.17- 23.50= 6.670C

Total volume=25.8+ 25.8=51.6ml

Density = mass/volume

Mass=51.6g

ΔH=51.6 ×4.184× 6.67=1440.02 J/g.0C